# Semantics-Empowered Space-Air-Ground-Sea Integrated Network: New Paradigm, Frameworks, and Challenges

### v1

@@ IV-C Techniques: Towards Meaning Enhancement
From a macro perspective, the meaning-oriented semantic communication system proposed in this section integrates the perception, communication, and computing parts. The coordinated collaboration of these three parts contributes to better application of the system in various scenarios within the SAGSIN. Now, let’s narrow our focus to some specific details and consider the practical works of system implementation. It is evident that DNNs serve as the core to build semantic codecs. However, considering the objective realities such as large-scale scenarios, highly dynamic channels, and sensitive communication security issues in the SAGSIN, it is necessary to incorporate additional targeted “designs” for the components of semantic communication systems. We name these “designs” as METs and categorize them into four major classes, which respectively focus on the channel adaptation, rate control, interpretability, and security. In the following, we will review the existing related works for each category, and the summary of these techniques is shown in Table IV.
@@ IV-C1 Channel-related METs
Analyzing and extracting the channel semantics to assist source information encoding is a new research direction [133, 134]. In [135], the authors introduces the concept of “channel semantics”, which encompasses parameter semantics (e.g., angle of departure, angle of arrival, number of paths, and so on) and environmental semantics (e.g., the layout, shape, and category of objects in the images). The system obtains environmental information in the SAGSIN through devices such as cameras and radars, and extracts channel semantics by DL techniques. The system considering channel semantics exhibits superiority in terms of meaning similarity metrics such as WER and FDSD. To address the impact of noise in complex channels, a confidence-based distillation mechanism is established in[88] to achieve the “joint semantic-noise coding (JSNC)”. The distillation time in the semantic distillation mechanism of JSNC can be automatically adjusted to accommodate the fluctuating channel conditions. Taking the example of the AWGN channel with 10 dB SNR, JSNC achieves a 24% lower WER compared to the Transformer baseline. Furthermore, considering the unique “semantic channel” (the information flow from semantic representation to semantic recovery) of the semantic communication system, the authors in [136] classify semantic noise for text sources into “literal semantic noise” and “adversarial semantic noise”, and propose a robust DL-based semantic communication system (R-DeepSC) to resist them. Specifically, a calibrated self-attention mechanism is utilized to counteract the former, and an adversarial training method is employed to combat the latter. Compared to other baselines considering physical noise only, R-DeepSC achieves superior BLEU and BERT scores at different SNRs. Effectively capturing and utilizing the rich CSI in the SAGSIN will bring extra recovery performance gains. The authors of [137] and [138] both design a plugin-in CSI modulation module, called “Channel ModNet”. [137] inserts it into the DeepJSCC decoder and proposes an adaptive semantic communication (ASC) based on the overfitting of the source and channel. This design adapts well to different SNRs in block fading channels. Reaching the same PSNR, ASC saves bandwidth up to 41% compared to the benchmark. In [138], the authors propose a wireless image transmission Transformer (WITT) model, utilizing the Swin Transformer and incorporating “Channel ModNet” into the codec. Compared to two baselines, namely the CNN-based DeepJSCC and the traditional BPG-LDPC SSCC, WITT exhibits superior PSNR and MS-SSIM at a wide range of (particularly low) SNRs. [139] proposes an improved version of the classical DeepJSCC, called the adaptive DeepJSCC. Specifically, an “attention feature (AF) module” is utilized to interact with codec DNNs, which can process the input CSI to adapt to different channel conditions. Experimental results show that the adaptive DeepJSCC achieves better PSNR than the classical DeepJSCC under a wide range of SNRs. In [140], the authors introduce a “response network” enabled nonlinear transform source-channel coding (NTSCC) framework to explicitly construct response functions and directly embed CSI into the parameters of the DeepJSCC codec. Under the AWGN channel with 10 dB SNR, this architecture achieves higher PSNR than traditional transmission schemes with standardized image codecs across all channel bandwidth ratios. Besides, a joint coding-modulation (JCM) scheme based on BPSK modulation for digital semantic communications is proposed in [141], matching the encoding and modulation processes with CSI better and achieving larger PSNR. Also, the lite distributed DL-based semantic communication system (L-DeepSC) in [142] is assisted by CSI during training, reducing the impact of fading channels on the transmission and increasing BLEU scores.
@@ IV-C2 Rate-control-related METs
Rate control is another measure to address the issues of highly dynamic channels for the SAGSIN. In the JSCC encoding process, allocating flexible rates to the extracted semantics allows for better adaptation to time-variant channels. Specifically, based on whether the retransmission protocols are adopted to prompt the correct decoding of the current codeword after decoding failure, rate-control techniques can be classified into retransmission-based ones and non-retransmission-based ones. A typical retransmission-based technique is the HARQ. Recently, a few variants of HARQ have been proposed in semantic communication systems. It is worth noting that the HARQs here are improved versions to adapt to the meaning-oriented semantic communication system, which is the main difference compared to the ones discussed in Section III-B. In [130] and [80], HARQ-IR used in traditional communication systems is improved to apply to semantic communication systems. The authors in [130] first propose SC-RS-HARQ, which belongs to the first category of DeepSSCC mentioned in Remark 4. The semantic codec is based on Transformers, while the channel codec utilize Reed-Solomon (RS) codes. Because HARQ-IR is only based on RS codes, the authors directly incorporate it into SC-RS-HARQ. To further improve system performance, they improve HARQ-IR and apply it to the DeepJSCC architecture, resulting in SCHARQ. Simulation results show that in a wide range of (especially low) SNRs, SCHARQ achieves lower WER compared to SC-RS-HARQ. [80] proposes a semantic video conferencing (SVC) network and improves HARQ-IR specifically for SVC, resulting in SVC-HARQ. The semantic error detector is employed to ascertain whether an incremental transmission is necessary for the received frame. Compared to traditional communication systems, SVC-HARQ exhibits lower AKD at high BERs. In [143], the authors propose a progressive semantic HARQ scheme based on the incremental knowledge (IK-HARQ). An encoding mechanism with adaptive semantic rate control is employed to dynamically adjust rates based on the contextual semantic complexity and channel conditions. Compared to other benchmarks, IK-HARQ outperforms in terms of BLEU in a wide range of SNRs. It is worth noting that non-retransmission does not imply non-feedback. In [144], no retransmission mechanism is employed, but the authors utilize feedback signals from the current codeword to guide the encoding of subsequent codewords, and introduce an autoencoder-based JSCC scheme, called DeepJSCC-f. They employ layered autoencoders and leverage channel output feedback to achieve variable codeword length transmission. The simulation results demonstrate the benefits of feedback in improving system performance. Compared to the non-feedback DeepJSCC and traditional communication systems, DeepJSCC-f reaches higher PSNR at different SNRs. The following works on rate control are all without retransmission or feedback. [145] proposes an adaptive JSCC for wireless image transmission. The “policy network” is introduced to activate or freeze visual features by dynamically generating binary masks, thereby achieving rate control. In scenarios with a high SNR and lower information content in the image, this model can learn reasonable strategies with less bandwidth. [146] and [147] propose the same “rate adaptive transmission” mechanism to apply to embedding vectors in latent representation of the encoder side. This enables the model to learn to allocate limited bandwidth resources in order to maximize overall performance. The model in [146] exhibits superior performance in terms of PSNR and MS-SSIM compared to DeepJSCC and traditional communication systems at various SNRs. In [147], the proposed architecture achieves up to 50% savings in channel bandwidth costs compared to traditional communication systems while maintaining the same PSNR and MS-SSIM performances. Besides, [148] introduces a variable-length semantic-channel coding (VL-SCC) method, where a “rate allocation network (RAN)” is utilized to estimate the optimal code length and enhancing the PSNR and LPIPS performance. The forward-adaption-based (FA-based) autoencoder scheme proposed in [127] improves performance in MS-SSIM and PSNR by introducing an “adaptive density learning module” at the transmitter to adaptively construct the codebook. [126] proposes a reinforcement learning based adaptive semantic coding (RL-ASC) model, which use a “rate controller” to adaptively adjust the code length. Under different rates, RL-ASC exhibits superior performance to benchmarks in terms of PSNR, SSIM, FID, and KID.
@@ IV-C3 Interpretability-related METs
One of the widely criticized flaws of DL is lack of interpretability. The black-box nature makes it hard for people to understand the internal decision process of models, hindering model improvement and optimization. Therefore, enhancing the interpretability of semantic communication systems by specific means (e.g., knowledge graph (KG) and probabilistic logic programming language (ProbLog)) can help further improve communication performance in the SAGSIN. Meanwhile, enhancing interpretability also facilitates the efficient learning of NNs by accelerating the convergence of complex models. This can better address the characteristic of limited device capabilities in the SAGSIN. KG is a graphical structure used to describe the relationships between entities. It is an extension and development of a knowledge base (KB). KGs contain a series of semantic triplets in the form of “entity-relation-entity.” As an advantage compared to KBs, KGs provide richer semantic associations and contextual information. Therefore, with the backdrop of building the meaning-oriented semantic communication system, KGs gain increasing attention recently. [149] proposes an interpretable semantic information detection algorithm by using KG triplets as semantic symbols. Also, an efficient semantic correction algorithm is also introduced by extracting inference rules from the KG. The KG-based semantic communication system in [149] exhibits high sentence similarity in a wide range of (particularly low) SNRs. The authors in [150] present a KG-enhanced semantic communication framework. In particular, they design a Transformer-based knowledge extractor in the decoder side to extract relevant factual triples based on the received noisy signal, thereby enhancing the semantic decoding capability. The experimental results demonstrate that regardless of channel types, in low SNRs, the semantic communication system based on knowledge extractor always yields a BLEU improvement of over 5% for the model. An interpretable semantic communication framework for text data transmission is considered in [73]. The semantic information is represented through a collection of semantic triples based on a KG. In addition, the authors also define a new metric, namely MSS, which comprehensively measures the accuracy and completeness of the reconstructed data. The framework proposed in [73], compared to traditional communication systems, can reduce the transmission data by 41.3% and improve the overall MSS by two times. ProbLog is a programming language that combines logic programming with probability theory, extending traditional logic programming. It represents knowledge in the form of logical clauses and allows specifying probability facts and rules within these logical clauses, enabling the representation and inference of knowledge involving uncertainty and probability. ProbLog has been widely applied in various fields, including machine learning (ML), NLP, and so on. In [151], the authors propose a semantic protocol model (SPM) to enhance the interpretability of the principles and impact of the NN-based protocol models (NPMs). Specifically, ProbLog is employed to convert the NPM into a symbolic graph. The experimental results confirm that the SPM performs closely to the NPM in performance while occupying only 0.02% of the memory. [152] introduces a KB-based on the ProbLog to facilitate semantic information exchange at the semantic communication level. In addition, [153] proposes semantic communications based on the “conceptual space.” Conceptual spaces are similar to KBs but abstract the semantic information measurement into a geometric space, defined as distances in the conceptual space. This approach also enhances the interpretability of the model.
@@ IV-C4 Security-related METs
Communication security, as a prerequisite for accurately transmitting information, has always been an essential concern, especially in the SAGSIN. This comprehensive network consists of numerous nodes, handling massive services at any given moment. When two specific nodes engage in the end-to-end communications, it is undesirable to be eavesdropped on the information by other unrelated nodes. Therefore, focusing on the security issues under the background of semantic communications becomes an emerging research direction [160]. A typical semantic communication system with eavesdropping is shown in Fig. 10. The first step of designing techniques to defend against attacks is understanding the attack mechanisms. A few works validate the impact of certain specific attacks on the performance of the semantic communication system, demonstrating the necessity of researching semantic communication security. [154] proposes a model called SemBLK, which can learn to generate destructive physical layer semantic attacks for end-to-end semantic communication systems in a black-box setting. Semantic perturbations are generated by introducing a surrogate semantic encoder in the SemBLK. Experimental results demonstrate the destructive and imperceptible nature of black-box attacks in the SemBLK by comparing metrics such as PSNR and SSIM with benchmarks. Similarly, in the recent work [155], a semantic perturbation generator called SemAdv is trained for physical-layer attacks that can pollute the images with specific semantics in an imperceptible, controllable, and input-agnostic manner. Besides, the authors in [156] demonstrate that the meaning-oriented semantic communication system exhibits weak resistance against backdoor (Trojan) attacks. The experimental results indicate that backdoor attacks are stealthy and selective. The second step is to design defense mechanisms to enhance the semantic security. [157] proposes a counter-eavesdropping DeepJSCC model, called DeepJSCEC. It can resist the eavesdroppers’ chosen-plaintext attacks without assuming their channel conditions. Compared to the SSCC encryption schemes of traditional communication systems, this model exhibits comparable or even superior PSNR, SSIM, and MS-SSIM performance at different SNRs. An encrypted semantic communication system (ESCS) is proposed in [131]. Specifically, the authors introduce an adversarial encryption training scheme, which enables the model to achieve high communication accuracy in both encrypted and unencrypted modes. Under the support of ESCS, at high SNRs, the BLEU score of Bob (receiver) approaches 1, while that of Eve (attacker) is less than 0.2. [158] proposes a StyleGAN-based semantic communication framework, which utilizes a privacy filter and a KB to replace private information with natural features in the KB, ensuring communication security. Compared to the benchmark (a kind of DeepJSCC), the model has a smaller and more stable LPIPS value under different SNRs. Some new security-based metrics used in semantic communications are also proposed in the literature. In [159], the authors examine and contrast traditional security methods, such as physical layer security, covert communications, and encryption, in terms of semantic information security. They introduce two metrics, which are semantic secrecy outage probability (SOP) for physical layer security techniques, and detection failure probability (DFP) for covert communication techniques. With these metrics as the ultimate goal, the security-related design becomes clearer.

### v3

@@ IV-C Techniques: Towards Meaning Enhancement
From a macro perspective, the meaning-oriented semantic communication system proposed in this section integrates the perception, communication, and computing parts. The coordinated collaboration of these three parts contributes to better application of the system in various scenarios within the SAGSIN. Now, let’s narrow our focus to some specific details and consider the practical works of system implementation. It is evident that DNNs serve as the core to build semantic codecs. However, considering the objective realities such as limited device capabilities, highly dynamic channels, and sensitive communication security issues in the SAGSIN, it is necessary to incorporate additional targeted “designs” for the components of semantic communication systems. We name these “designs” as METs and categorize them into four major classes, which respectively focus on the channel adaptation, rate control, interpretability, and security. In the following, we will review the existing related works for each category, and the summary of these techniques is shown in Table VII.
@@ IV-C1 Channel-related METs
Analyzing and extracting the channel semantics to assist source information encoding is a new research direction [147, 148]. In [149], the authors introduce the concept of “channel semantics”, which encompasses parameter semantics (e.g., angle of departure, angle of arrival, number of paths, and so on) and environmental semantics (e.g., the layout, shape, and category of objects in the images). The system obtains environmental information in the SAGSIN through devices such as cameras and radars, and extracts channel semantics by DL techniques. The system considering channel semantics exhibits superiority in terms of meaning similarity metrics such as WER and FDSD.
To address the impact of noise in complex channels, a confidence-based distillation mechanism is established in[102] to achieve the “joint semantic-noise coding (JSNC)”. The distillation time in the semantic distillation mechanism of JSNC can be automatically adjusted to accommodate the fluctuating channel conditions. Taking the example of the AWGN channel with 10 dB SNR, JSNC achieves a 24% lower WER compared to the Transformer baseline. Furthermore, considering the unique “semantic channel” (the information flow from semantic representation to semantic recovery) of the semantic communication system, the authors in [150] classify semantic noise for text sources into “literal semantic noise” and “adversarial semantic noise”, and propose a robust DL-based semantic communication system (R-DeepSC) to resist them. Specifically, a calibrated self-attention mechanism is utilized to counteract the former, and an adversarial training method is employed to combat the latter. Compared to other baselines considering physical noise only, R-DeepSC achieves superior BLEU and BERT scores at different SNRs.
Effectively capturing and utilizing the rich CSI in the SAGSIN will bring extra recovery performance gains. The authors of [151] design a plugin-in CSI modulation module, called “Channel ModNet”. [151] inserts it into the DeepJSCC decoder and proposes an adaptive semantic communication (ASC) based on the overfitting of the source and channel. Different from offline learning method adopted by most works on semantic communications, [151] firstly introduces online learning policies, where the model parameters will be updated in model inference procedure besides training. This design adapts well to different SNRs in block fading channels. Reaching the same PSNR, ASC saves bandwidth up to 41% compared to the benchmark. Similarly, [153] also adopts an online design in order to adapt to dynamic channel conditions. The authors introduce a “response network” enabled nonlinear transform source-channel coding (NTSCC) framework to explicitly construct response functions and directly embed CSI into the parameters of the DeepJSCC codec. Under the AWGN channel with 10 dB SNR, this architecture achieves higher PSNR than traditional transmission schemes with standardized image codecs across a wide range of channel bandwidth ratios.
@@
Also, in [152], the authors propose a wireless image transmission Transformer (WITT) model, utilizing the Swin Transformer and incorporating “Channel ModNet”[151] into the codec. Compared to two baselines, namely the CNN-based DeepJSCC and the traditional BPG-LDPC SSCC, WITT exhibits superior PSNR and MS-SSIM at a wide range of (particularly low) SNRs. [154] proposes an improved version of the classical DeepJSCC, called the adaptive DeepJSCC. Specifically, an “attention feature module” is utilized to interact with codec DNNs, which can process the input CSI to adapt to different channel conditions. Experimental results show that the adaptive DeepJSCC achieves better PSNR than the classical DeepJSCC under a wide range of SNRs. Besides, a joint coding-modulation scheme based on binary phase shift keying (BPSK) modulation for digital semantic communications is proposed in [155], matching the encoding and modulation processes with CSI better and achieving larger PSNR. Also, a lite distributed DL-based semantic communication system in [156] is assisted by CSI during training, reducing the impact of fading channels on the transmission and increasing BLEU scores.
Remark 6.
Though introducing online learning method will achieve performance gain owing to adaptability to dynamic channels[151, 153], the inherent spacial/temporal computing complexity cannot be neglected, especially for the devices with limited capabilities distributed in SAGSIN. Specifically, the deep networks utilized in semantic extraction & reconstruction possess huge amount of parameters, and updating such networks will consume caching resources and cause long processing delays. Therefore, striking a trade-off between performance gain and computing complexity in online learning policy will be a promising idea.
@@ IV-C2 Rate-control-related METs
Rate control is another measure to address the issues of highly dynamic channels for the SAGSIN. In the JSCC encoding process, allocating flexible rates to the extracted semantics allows for better adaptation to time-variant channels. Specifically, based on whether the retransmission protocols are adopted to prompt the correct decoding of the current codeword after decoding failure, rate-control techniques can be classified into retransmission-based ones and non-retransmission-based ones.
@@
A typical retransmission-based technique is the HARQ. Recently, a few variants of HARQ have been proposed in semantic communication systems. It is worth noting that the HARQs here are improved versions to adapt to the meaning-oriented semantic communication system, which is the main difference compared to the ones discussed in Section III-B. In [144] and [94], HARQ-IR used in traditional communication systems is improved to apply to semantic communication systems. The authors in [144] firstly propose SC-RS-HARQ, which belongs to the first category of DeepSSCC mentioned in Remark 5. The semantic codec is based on Transformers, while the channel codec utilize Reed-Solomon (RS) codes. Because HARQ-IR is only based on RS codes, the authors directly incorporate it into SC-RS-HARQ. To further improve system performance, they improve HARQ-IR and apply it to the DeepJSCC architecture, resulting in SCHARQ. Simulation results show that in a wide range of (especially low) SNRs, SCHARQ achieves lower WER compared to SC-RS-HARQ. [94] proposes a semantic video conferencing (SVC) network and improves HARQ-IR specifically for SVC, resulting in SVC-HARQ. The semantic error detector is employed to ascertain whether an incremental transmission is necessary for the received frame. Compared to traditional communication systems, SVC-HARQ exhibits lower AKD at high BERs. In [157], the authors propose a progressive semantic HARQ scheme based on the incremental knowledge (IK-HARQ). An encoding mechanism with adaptive semantic rate control is employed to dynamically adjust rates based on the contextual semantic complexity and channel conditions. Compared to other benchmarks, IK-HARQ outperforms in terms of BLEU in a wide range of SNRs.
It is worth noting that non-retransmission does not imply non-feedback. In [158], no retransmission mechanism is employed, but the authors utilize feedback signals from the current codeword to guide the encoding of subsequent codewords, and introduce an autoencoder-based JSCC scheme, called DeepJSCC-f. They employ layered autoencoders and leverage channel output feedback to achieve variable codeword length transmission. The simulation results demonstrate the benefits of feedback in improving system performance. Compared to the non-feedback DeepJSCC and traditional communication systems, DeepJSCC-f reaches higher PSNR at different SNRs.
The following works on rate control are all without retransmission or feedback. [159] proposes an adaptive JSCC for wireless image transmission. The “policy network” is introduced to activate or freeze visual features by dynamically generating binary masks, thereby achieving rate control. In scenarios with a high SNR and lower information content in the image, this model can learn reasonable strategies with less bandwidth. [160] and [161] propose the same “rate adaptive transmission” mechanism to apply to embedding vectors in latent representation of the encoder side. This enables the model to learn to allocate limited bandwidth resources in order to maximize overall performance. The model in [160] exhibits superior performance in terms of PSNR and MS-SSIM compared to DeepJSCC and traditional communication systems at various SNRs. In [161], the proposed architecture achieves up to 50% savings in channel bandwidth costs compared to traditional communication systems while maintaining the same PSNR and MS-SSIM performances. Besides, [162] introduces a variable-length semantic-channel coding method, where a “rate allocation network” is utilized to estimate the optimal code length and enhance the PSNR and LPIPS performance. The forward-adaption-based autoencoder scheme proposed in [141] improves performance in MS-SSIM and PSNR by introducing an “adaptive density learning module” at the transmitter to adaptively construct the codebook. [140] proposes a reinforcement learning based adaptive semantic coding (RL-ASC) model, which uses a “rate controller” to adaptively adjust the code length. Under different rates, RL-ASC exhibits superior performance to benchmarks in terms of PSNR, SSIM, FID, and KID.
@@ IV-C3 Interpretability-related METs
One of the widely criticized flaws of DL is lack of interpretability. The black-box nature makes it hard for people to understand the internal decision process of models, hindering model improvement and optimization. Therefore, enhancing the interpretability of semantic communication systems by specific means (e.g., KG and probabilistic logic programming language (ProbLog)) can help further improve communication performance in the SAGSIN. Meanwhile, enhancing interpretability also facilitates the efficient learning of NNs by accelerating the convergence of complex models. This can better address the characteristic of limited device capabilities in the SAGSIN.
@@
KG is a graphical structure used to describe the relationships between entities. It is an extension and development of a knowledge base (KB). KGs contain a series of semantic triplets in the form of “entity-relation-entity”. As an advantage compared to KBs, KGs provide richer semantic associations and contextual information. Therefore, with the backdrop of building the meaning-oriented semantic communication system, KGs gain increasing attention recently. [163] proposes an interpretable semantic information detection algorithm by using KG triplets as semantic symbols. Also, an efficient semantic correction algorithm is also introduced by extracting inference rules from the KG. The KG-based semantic communication system in [163] exhibits high sentence similarity in a wide range of (particularly low) SNRs. The authors in [164] present a KG-enhanced semantic communication framework. In particular, they design a Transformer-based knowledge extractor in the decoder side to extract relevant factual triplets based on the received noisy signal, thereby enhancing the semantic decoding capability. The experimental results demonstrate that regardless of channel types, in low SNRs, the semantic communication system based on knowledge extractor always yields a BLEU improvement of over 5% for the model. An interpretable semantic communication framework for text data transmission is considered in [87]. The semantic information is represented through a collection of semantic triplets based on a KG. In addition, the authors also define a new metric, namely MSS, which comprehensively measures the accuracy and completeness of the reconstructed data. The framework proposed in [87], compared to traditional communication systems, can reduce the amount of transmission data by 41.3% and improve the overall MSS by two times.
ProbLog is a programming language that combines logic programming with probability theory, extending traditional logic programming. It represents knowledge in the form of logical clauses and allows specifying probability facts and rules within these logical clauses, enabling the representation and inference of knowledge involving uncertainty and probability. ProbLog has been widely applied in various fields, including machine learning (ML), NLP, and so on. In [165], the authors propose a semantic protocol model (SPM) to enhance the interpretability of the NN-based protocol models (NPMs). Specifically, ProbLog is employed to convert the NPM into a symbolic graph. The experimental results confirm that the SPM performs closely to the NPM while occupying only 0.02% of the memory. [166] introduces a KB-based model interpreted by the ProbLog to facilitate semantic information exchange at the semantic communication level.
In addition, [167] proposes semantic communications based on the conceptual space. Conceptual spaces are similar to KBs, wherein the coder maps the semantic information to the coordinates of a geometric space. By this means, semantic distortion is quantified as the distance between received information and original information in the conceptual space. This approach also enhances the interpretability of the model.
@@ IV-C4 Security-related METs
Communication security, as a prerequisite for accurately transmitting information, has always been an essential concern, especially in the SAGSIN. This comprehensive network consists of numerous nodes, handling massive services at any given moment. When two specific nodes engage in the end-to-end communications, they do not want to be eavesdropped by other unrelated nodes. Therefore, focusing on the security issues under the background of semantic communications becomes an emerging research direction [174]. A typical semantic communication system with eavesdropping is shown in Fig. 10.
The first step of designing techniques to defend against attacks is understanding the attack mechanisms. A few works validate the impact of certain specific attacks on the performance of the semantic communication system, demonstrating the necessity of researching semantic communication security. [168] proposes a model called SemBLK, which can learn to generate destructive physical layer semantic attacks for end-to-end semantic communication systems in a black-box setting. Semantic perturbations are generated by introducing a surrogate semantic encoder in the SemBLK. Experimental results demonstrate the destructive and imperceptible nature of black-box attacks in the SemBLK by comparing metrics such as PSNR and SSIM with benchmarks. Similarly, in the recent work [169], a semantic perturbation generator called SemAdv is trained for physical-layer attacks that can pollute the images with specific semantics in an imperceptible, controllable, and input-agnostic manner. Besides, the authors in [170] demonstrate that the meaning-oriented semantic communication system exhibits weak resistance against backdoor (Trojan) attacks. The experimental results indicate that backdoor attacks are stealthy and selective.
@@
The second step is to design defense mechanisms to enhance the semantic security. [171] proposes a counter-eavesdropping DeepJSCC model, called DeepJSCEC. It can resist the eavesdroppers’ chosen-plaintext attacks without assuming their channel conditions. Compared to the SSCC encryption schemes of traditional communication systems, this model exhibits comparable or even superior PSNR, SSIM, and MS-SSIM performance at different SNRs. An encrypted semantic communication system (ESCS) is proposed in [145]. Specifically, the authors introduce an adversarial encryption training scheme, which enables the model to achieve high communication accuracy in both encrypted and unencrypted modes. Under the support of ESCS, at high SNRs, the BLEU score of Bob (receiver) approaches 1, while that of Eve (attacker) is less than 0.2. [172] proposes a StyleGAN-based semantic communication framework, which utilizes a privacy filter and a KB to replace private information with natural features in the KB, ensuring communication security. Compared to the benchmark (a kind of DeepJSCC), the model has a smaller and more stable LPIPS value under different SNRs.
Some new security-based metrics used in semantic communications are also proposed in the literature. In [173], the authors examine and contrast traditional security methods, such as physical layer security, covert communications, and encryption, in terms of semantic information security. They introduce two metrics, which are semantic secrecy outage probability (SOP) for physical layer security techniques, and detection failure probability (DFP) for covert communication techniques. With these metrics as the ultimate goal, the security-related design becomes clearer.